# Relation Also Knows: Rethinking the Recall and Editing of Factual Associations in Auto-Regressive Transformer Language Models

### v1

@@ Abstract
The storage and recall of factual associations in auto-regressive transformer language models (LMs) have drawn a great deal of attention, inspiring knowledge editing by directly modifying the located model weights. Most editing works achieve knowledge editing under the guidance of existing interpretations of knowledge recall that mainly focus on subject knowledge. However, these interpretations are seriously flawed, neglecting relation information and leading to the over-generalizing problem for editing. In this work, we discover a novel relation-focused perspective to interpret the knowledge recall of transformer LMs during inference and apply it on knowledge editing to avoid over-generalizing. Experimental results on the dataset supplemented with a new R-Specificity criterion demonstrate that our editing approach significantly alleviates over-generalizing while remaining competitive on other criteria, breaking the domination of subject-focused editing for future research. [^1] The code and data will be released at Github upon publication.
@@ Introduction
Language models are often regarded as knowledge bases, storing factual associations in the parameters which can be simply recalled through prompting (Petroni et al. 2019; Lester, Al-Rfou, and Constant 2021; Jiang et al. 2020; Roberts, Raffel, and Shazeer 2020; Petroni et al. 2020; Heinzerling and Inui 2021; Wang, Liu, and Zhang 2021). For instance, for the factual association shown in triplet <Marco Reus, citizen-of, O> with the subject Marco Reus and the relation citizen-of, the object O can be obtained from the next token prediction of GPT-like language models given the prompt "Marco Reus is a citizen of". Recent works investigate where factual knowledge is stored and how the factual knowledge is extracted from auto-regressive transformer LMs, suggesting that the feedforward MLP sublayer performs as key-value memories which is the key component for the storing and recall of factual associations (Geva et al. 2021, 2022, 2023). The sight into the interpretation of auto-regressive transformer LMs makes renewing their knowledge by directly modifying the MLP weights possible, inspiring knowledge editing via the locate-then-edit paradigm that modifies the located weights (Yao et al. 2023; Meng et al. 2022a, b; Li et al. 2024a). This sort of methods provide convenience for altering the behavior of models permanently on a small amount of facts while ensuring least side-effects, especially meaningful in the era of large language models.
@@
We conjecture that these locate-then-edit methods suffer from over-generalizing since they only focus on subjects and fail to take relations of factual associations into consideration during editing. Thus we firstly investigate what happens on relation tokens (e.g. "was born in") in knowledge recall during inference to understand why previous works fail to take relations into account. The interpretation of knowledge recall involves which positions of tokens and which layers of auto-regressive transformer LMs primarily contribute to the prediction, and what interpretable information is encoded at the corresponding points. Through causal tracing (Meng et al. 2022a) for relations, we discover that the most contributing MLP and multi-head self attention MHSA sublayers for the propagation of relation representations appear at the last relation token. Furthermore, we analyze the trend of attributes rate and the target object ranking flow via the vocabulary lens (Geva et al. 2021) of hidden representations at the identified last relation token across layers. Through the analysis results, we conclude with the relation-focused interpretation of knowledge recall that relation-related attributes (i.e. relational knowledge) are aggregated from the first layer till middle-late layers at the last relation token and that the target object token is extracted from the aggregated relational knowledge. We also validates the importance of MLP over MHSA for the aggregation of relational knowledge by the decline of attributes rate after blocking the MLP and MHSA sublayers respectively during inference. According to the investigation results, we notice that the inference of relations takes place at the last relation token and practically completes in middle-late layers. However, previous works achieve editing via modification of MLP in the middle-early layer at the last subject token, earlier than the inference of relations is completed and also unable to attend to the last relation token behind due to the nature of auto-regressive transformer LMs. As a result, previous locate-then-edit methods fail to take relational knowledge into account and tend to modify no matter what relationships of the target editing subject, leading to the over-generalizing problem.
@@
In order to take relations into consideration, we propose to edit under the guidance of the novel relation-focused interpretation that modifies the MLP in the end of aggregation of relational knowledge (i.e. in the middle-late layer at the last relation token). Although simply editing at this point can attend to the subject, it loses the specificity on predictions of prompts with the same relation but different subjects with target editing (i.e. neighborhood subject prompts). Therefore, to make the hidden representations of neighborhood subject prompts more distinguishable at this point, we add an optimization target to the deduction of the weight modification to enhance the difference between such neighborhood prompts, constraining the editing to the certain subject. To sum up, we propose the Relation-focused Editing for auto-regressive Transformer LMs with Subject constraints (RETS) method to solve the problem of over-generalizing in knowledge editing and verify the reliability of the novel relation-focused interpretation.
@@
For evaluation, we supplement the COUNTERFACT  (Meng et al. 2022a) dataset with a new criterion Relation Specificity (i.e. R-Specificity) that measures the influence on unrelated facts of the edited subject. Experimental results on the supplemented dataset show that our editing method avoids over-generalizing by outperforming the state-of-the-art locate-then-edit methods over $30\%$ on Relation Specificity, while remaining competitive with the baselines on the previous criteria. Our strategy of knowledge editing exhibits the most balanced performance overall and also validates the relation-focused interpretation on the recall of factual associations in auto-regressive transformer LMs, providing a novel perspective for future research on knowledge editing and the recall mechanism.
@@
Knowledge editing methods intend to alter the behavior of language models within the domain related to the edited fact, avoiding side-effects on unrelated facts (Yao et al. 2023; Dai et al. 2022; De Cao, Aziz, and Titov 2021; Dong et al. 2022; Mitchell et al. 2022; Hase et al. 2023). A line of locate-then-edit methods are proposed with the support of the recall mechanisms mentioned above, localizing a decisive MLP weight in middle-early layers at the last-subject position and directly modify it through rank-one model editing ROME (Meng et al. 2022a) for each single factual association. MEMIT (Meng et al. 2022b) improve ROME to be applicable on numerous edits simultaneously by spreading the updates evenly over several middle MLP sublayers while processing the subject representation. PMET (Li et al. 2024a) further obtains more precise FFN output at the last-subject position for editing by taking both MHSA and FFN information into consideration during optimization.
@@
(a) Average attributes rate
(b) Attributes rate decline
(c) Average token ranking
Figure 2: The factual information detection on the vocabulary lens of the last-relation representation for GPT2-XL. (a) The average attributes rates of the representations as shown in yellow bars. (b) The average attributes rate decline at 48-th layer while blocking the MLP or MHSA sublayer respectively. (c) The average rankings of the target objects and random tokens.
@@ Results.
Figure 3 shows the average indirect effect of relation (AIER) heatmaps on MLP and MHSA sublayers for GPT2-XL. We note that for both MLP and MHSA, the most significant output representations are detected during inference at the last-relation token (also the last input token). Surprisingly, the AIER of relation prefixes ("rp" in Figure 3) contributes little to the prediction. All that matters for the entire relation information is the hidden representation at the last-relation token. Appendix A shows the AIER heatmaps on some specific cases. For quantitative comparisons, the average AIER across layers for tokens of certain positions is shown in Table 1. The average AIER at the last-relation position is far beyond that at any other position. This indicates that the MLP and MHSA sublayers are most active at the last-relation token which is the decisive position that process the potential factual information of the relation representation. Moreover, we can tell from the heatmaps (Figure 3) that active MLP sublayers amass from the early layers to the middle-late site of the model, indicating that the update process for the relation representation is of long duration and finishes at the middle-late layers. Therefore, we conclude that the inference of relations primarily takes place at the last relation token from early layers to middle-late layers.
(a) AIER heatmap of MLP
(b) AIER heatmap of MHSA
Figure 3: Average Indirect Effect of Relation results for MLP and MHSA sublayers over 1000 facts on GPT2-XL. "rp" stands for the relation prefix in front of the subject (e.g. "The mother tongue language of" in the input prompt "The mother tongue language of Isabelle Breitman is"). "fs", "ms" and "ls" stand for the first-subject token, middle-subject tokens and last-subject token. "fr", "mr" and "lr" stand for the first-relation token, middle-relation tokens and last-relation token. "*" marks the intervened tokens in the corrupted run.
@@
Figure 2(c) plots the average rankings of the target objects and random tokens in the vocabulary distributions induced at the last-relation position. We can tell from the line charts that the average rankings of the target objects has been rising from early layers to middle-late layers, while the average ranking for random tokens remains low in all layers in comparison. This indicates the target objects are promoted to the final prediction gradually since the first layer of the models. Table 2 shows the proportion of the 1000 facts where the correctly predicted objects are included in corresponding $\mathbb{A}_{relation}(r)$ and the Spearman rank correlation coefficient between the average negative rankings of the objects and the average attributes rate of the representations of the last-relation position across layers. For GPT2-XL (GPT-J), $68\%$ ($83\%$) of correctly predicted objects are included in the corresponding $\mathbb{A}_{relation}(r)$ and the Spearman rank coefficient is $0.97$ ($0.73$), a strongly positive correlation between the extraction of the target objects and the accumulation of relation-related attributes. Thus, we conclude with the relation-focused interpretation that target objects are extracted from the relation-related attributes which are enriched at the last-relation token from early layers till middle-late layers and the MLP sublayers are essential in the update of relation representations. Therefore, under the guidance of this interpretation, we achieve editing by modifying the MLP sublayer in end of aggregation of relational knowledge (i.e. in the middle-late layer) with the relation representation (i.e. at the last-relation token) while taking subjects into account.
Knowledge Editing from Relation-focused Perspective
@@
(a) Efficacy
(b) Generalization
(c) S-Specificity
(d) R-Specificity
(e) Score
Figure 5: The performance of RETS (purple lines) editing at the last-relation token on different layers compared with ROME (orange line) editing at the last-subject token for 50 prompts. Std deviation is shown in areas.
@@
Our editing experiments are mainly conducted on GPT2-XL and GPT-J. We also evaluate the basic editing performance on Llama-2 (7B). To compare with, we choose the methods for editing each single factual association as our baselines, including Constrained Fine-Tuning (FT+L) (Zhu et al. 2020), the meta-learning method (MEND) (Mitchell et al. 2022) which learns the update of model weights with additional networks, ROME (Meng et al. 2022a) on the subject representation in the early site and the improved precise model editing (PMET) (Li et al. 2024a) which optimizes the parallel MHSA and MLP representations simultaneously.
For evaluation, we conduct experiments on $10,000$ ($2,000$) samples of COUNTERFACT (Meng et al. 2022a) dataset for GPT2-XL (GPT-J), $1,000$ samples of COUNTERFACT for Llama-2 and $10,000$ samples of Zero-shot Relation Extraction (zsRE) (Levy et al. 2017; Meng et al. 2022a) for GPT2-XL. We edit on the $36$-th layer of GPT2-XL, $18$-th layer of GPT-J and $23$-th layer of Llama-2. For COUNTERFACT, we supplement this dataset with unrelated facts of the same subject for each target editing and the corresponding metric Relation Specificity (R-Specificity) to measure the over-generalizing problem. To be specific, given the edited fact <$s_{1},r_{1},o_{1}^{*}$> and the unrelated fact with the same subject <$s_{1},r_{2},o_{2}$>, we test $\mathbb{P}[o_{2}]>\mathbb{P}[o_{1}^{*}]$ as the R-Specificity Score with the input prompt of <$s_{1},r_{2}$>, computed similarly to the existing metrics. We also test the original metrics including the basic Efficacy accuracy Scores to measure the success rate of target editing with $\mathbb{P}[o_{1}^{*}]>\mathbb{P}[o_{1}]$ where $o_{1}$ is the original object of <$s_{1},r_{1}$>, the Generalization accuracy score for generalization on the paraphrased statement <$s_{1},r_{1}^{\prime}$> with $\mathbb{P}[o_{1}^{*}]>\mathbb{P}[o_{1}]$, the renamed Subject Specificity S-Specificity accuracy score for specificity within neighborhood subjects <$s_{3},r_{1}$> with the original object $o_{3}$ that $\mathbb{P}[o_{1}^{*}]<\mathbb{P}[o_{3}]$. The advanced Fluency and Consistency scores measure the quality of generated full texts. Higher scores indicate better performance for all metrics. For zsRE, the metics and evaluation results are displayed in Appendix G. Details for the construction of supplemented samples on R-Specificity and detailed settings for implementations are presented in Appendix F.
Table 3: The evaluation results on COUNERTFACT for GPT2-XL and GPT-J. The significantly failed values for the editing methods on basic criteria are presented in red. "Score" shows the average value on the basic criteria: Efficacy, Generalization, S-Specificity and R-Specificity. "SC" stands for the subject constraints on our relation-focused editing. R-Specificity values for raw models are $100.0\%$ since the criterion is constructed according to the top token predictions of the raw models.
@@
The ablation of the subject constraints for editing COUNTERFACT leads to the $35\%$ decrease on Entity Specificity for GPT2-XL, which indicates the effectiveness of the subject constraints for the relation-focused editing. Despite of the results on Generalization and Entity Specificity where RETS loses about $20\%$ and $10\%$ respectively compared with the subject-centered editing methods, RETS exhibits the most balanced performance with its simple way of combining the subject information into editing, which shows the potential of editing from relation-focused perspective.
We note that the quality of generated texts denoted by Fluency and Consistency for RETS declines due to the subject constraints that disrupt the associative memories of the original $K$ and $V$ in Eqn. 13, which is verified by the ablation study where Fluency increases sharply without the subject constraints. It calls for better ways of combining subject constraints into editing but does not detract from the novelty of our interpretation and editing from the relation-focused perspective.
Figure 6: An anecdotal example of the correct behavior for RETS and the incorrect behavior for ROME on GPT2-XL. Predictions in red denote unexpectedly changed answers.
@@
We further test the effectiveness of RETS while editing on different layers and compare it with the behavior of ROME which edits at the last-subject token . Figure 5 plots the editing performance on four criteria (a,b,c,d) and the average scores (e) across layers. The performance of RETS vibrates before middle layers and become stable since middle-late layers, validating our interpretation of relation-focused knowledge recall that the object is attracted from relation-related attributes which are accumulated before middle-late layers. RETS editing on middle-late layers shows more balanced performance than editing on any layer at the last-subject token where the relation information behind can hardly propagates to.

### v2

@@ Abstract
The storage and recall of factual associations in auto-regressive transformer language models (LMs) have drawn a great deal of attention, inspiring knowledge editing by directly modifying the located model weights. Most editing works achieve knowledge editing under the guidance of existing interpretations of knowledge recall that mainly focus on subject knowledge. However, these interpretations are seriously flawed, neglecting relation information and leading to the over-generalizing problem for editing. In this work, we discover a novel relation-focused perspective to interpret the knowledge recall of transformer LMs during inference and apply it on single knowledge editing to avoid over-generalizing. Experimental results on the dataset supplemented with a new R-Specificity criterion demonstrate that our editing approach significantly alleviates over-generalizing while remaining competitive on other criteria, breaking the domination of subject-focused editing for future research.
@@ Introduction
Language models are often regarded as knowledge bases, storing factual associations in parameters which can be simply recalled through prompting (Petroni et al. 2019; Lester, Al-Rfou, and Constant 2021; Jiang et al. 2020; Roberts, Raffel, and Shazeer 2020; Petroni et al. 2020; Heinzerling and Inui 2021; Wang, Liu, and Zhang 2021). For instance, for the factual association shown in triplet <Marco Reus, citizen-of, O> with the subject Marco Reus and the relation citizen-of, the object O can be obtained from the next token prediction of GPT-like language models given the prompt "Marco Reus is a citizen of". Recent works investigate where factual knowledge is stored and how the factual knowledge is extracted from auto-regressive transformer LMs, suggesting that the feedforward MLP sublayer performs as key-value memories which is the key component for the storing and recall of factual associations (Geva et al. 2021, 2022, 2023).
@@
The sight into the interpretation of auto-regressive transformer LMs makes renewing their knowledge by directly modifying the MLP weights possible, inspiring knowledge editing via the locate-then-edit paradigm that modifies the located weights (Yao et al. 2023; Meng et al. 2022a, b; Li et al. 2024a). This sort of methods provide convenience for altering the behavior of models permanently on a small amount of facts while ensuring least side-effects, especially meaningful in the era of large language models.
@@
We conjecture that these locate-then-edit methods suffer from over-generalizing since they only focus on subjects and fail to take relations of factual associations into consideration during editing. Thus we firstly investigate what happens on relation tokens (e.g. "was born in") in knowledge recall during inference to understand why previous works fail to take relations into account. The interpretation of knowledge recall involves which positions of tokens and which layers of auto-regressive transformer LMs primarily contribute to the prediction, and what interpretable information is encoded at the corresponding points. Through causal tracing (Meng et al. 2022a) for relations, we discover that the most contributing MLP and multi-head self attention MHSA sublayers for the propagation of relation representations appear at the last relation token. Furthermore, we analyze the trend of attributes rate and the target object ranking flow via the vocabulary lens (Geva et al. 2021) of hidden representations at the identified last relation token across layers. Through the analysis results, we conclude with the relation-focused interpretation of knowledge recall that relation-related attributes (i.e. relational knowledge) are aggregated from the first layer till middle-late layers at the last relation token and that the target object token is extracted from the aggregated relational knowledge. We also validate the importance of MLP over MHSA for the aggregation of relational knowledge by the decline of attributes rate after blocking the MLP and MHSA sublayers respectively during inference. According to the investigation results, we notice that the inference of relations takes place at the last relation token and practically completes in middle-late layers. However, previous works achieve editing via modification of MLP in the middle-early layer at the last subject token, earlier than the inference of relations is completed and also unable to attend to the last relation token behind due to the nature of auto-regressive transformer LMs. As a result, previous locate-then-edit methods fail to take relational knowledge into account and tend to modify no matter what relationships of the target editing subject, leading to the over-generalizing problem.
@@
In order to take relations into consideration, we propose to edit under the guidance of the novel relation-focused interpretation that modifies the MLP in the end of aggregation of relational knowledge (i.e. in the middle-late layer at the last relation token). Although simply editing at this point can attend to the subject, it loses the specificity on predictions of prompts with the same relation but different subjects with target editing (i.e. neighborhood subject prompts). Therefore, to make the hidden representations of neighborhood subject prompts more distinguishable at this point, we add an optimization target to the deduction of the weight modification to enhance the difference between such neighborhood prompts, constraining the editing to the certain subject. To sum up, we propose the Relation-focused Editing for auto-regressive Transformer LMs with Subject constraints (RETS) method to solve the problem of over-generalizing in single knowledge editing initially and verify the reliability of the novel relation-focused interpretation.
@@
For evaluation, we supplement the COUNTERFACT  (Meng et al. 2022a) dataset with a new criterion Relation Specificity (i.e. R-Specificity) that measures the influence on unrelated facts of the edited subject. Experimental results on the supplemented dataset show that our editing method avoids over-generalizing by outperforming the state-of-the-art locate-then-edit methods over $30\%$ on Relation Specificity, while remaining competitive with the baselines on the previous criteria. Our strategy of single knowledge editing exhibits the most balanced performance overall and also validates the relation-focused interpretation on the recall of factual associations in auto-regressive transformer LMs, providing a novel perspective for future research on knowledge editing and the recall mechanism.
@@
Knowledge editing methods intend to alter the behavior of language models within the domain related to the edited fact, avoiding side-effects on unrelated facts (Yao et al. 2023; Dai et al. 2022; De Cao, Aziz, and Titov 2021; Dong et al. 2022; Mitchell et al. 2022; Hase et al. 2023). A line of locate-then-edit methods are proposed with the support of the recall mechanisms mentioned above, localizing a decisive MLP weight in middle-early layers at the last-subject position and directly modify it through rank-one model editing ROME (Meng et al. 2022a) for each single factual association. MEMIT (Meng et al. 2022b) improve ROME to be applicable on numerous edits simultaneously by spreading the update evenly over several middle MLP sublayers while processing the subject representation. PMET (Li et al. 2024a) further obtains more precise FFN output at the last-subject position for editing by taking both MHSA and FFN information into consideration during optimization.
@@
Figure 2: Average Indirect Effect of Relation results for MLP and MHSA sublayers over 1000 facts on GPT2-XL. X-axis shows the layers. "rp" stands for the relation prefix in front of the subject (e.g. "The mother tongue language of" in the input prompt "The mother tongue language of Isabelle Breitman is"). "fs", "ms" and "ls" stand for the first-subject token, middle-subject tokens and last-subject token. "fr", "mr" and "lr" stand for the first-relation token, middle-relation tokens and last-relation token. "*" marks the intervened tokens in the corrupted run.
@@
Figure 3: The factual information detection on the vocabulary lens of the last-relation representation for GPT2-XL over 1000 prompts. (a) The average attributes rates as shown in yellow bars. (b) The average attributes rate decline at 48-th layer while blocking the MLP or MHSA sublayer respectively. (c) The average rankings of the target objects and random tokens.
@@ Results.
Figure 2 shows the average indirect effect of relation (AIER) heatmaps on MLP and MHSA sublayers for GPT2-XL. We note that for both MLP and MHSA, the most significant output representations are detected during inference at the last-relation token (also the last input token). Surprisingly, the AIER of relation prefixes ("rp" in Figure 2) contributes little to the prediction. All that matters for the entire relation information is the hidden representation at the last-relation token. Appendix A shows the AIER heatmaps on some specific cases. For quantitative comparisons, the average AIER across layers for tokens of certain positions is shown in Table 1. The average AIER at the last-relation position is far beyond that at any other position. This indicates that the MLP and MHSA sublayers are most active at the last-relation token which is the decisive position that process the potential factual information of the relation representation. Moreover, we can tell from Figure 2(a) that active MLP sublayers amass from the early layers to the middle-late site of the model, indicating that the update process for the relation representation is of long duration and finishes at the middle-late layers. Therefore, we conclude that the inference of relations primarily takes place at the last relation token from early layers to middle-late layers.
@@
Figure 3(c) plots the average rankings of the target objects and random tokens in the vocabulary distributions induced at the last-relation position. We can tell from the line charts that the average rankings of the target objects has been rising from early layers to middle-late layers, while that of random tokens remains low in all layers in comparison. This indicates the target objects are promoted to the final prediction gradually since the first layer of the models. Table 2 shows the proportion of the 1000 facts where the correctly predicted objects are included in corresponding $\mathbb{A}_{relation}(r)$ and the Spearman rank correlation coefficient between the average negative rankings of the objects and the average attributes rate of the representations of the last-relation position across layers. For GPT2-XL (GPT-J), $68\%$ ($83\%$) of correctly predicted objects are included in the corresponding $\mathbb{A}_{relation}(r)$ and the Spearman rank coefficient is $0.97$ ($0.73$), a strongly positive correlation between the extraction of the target objects and the accumulation of relation-related attributes. Thus, we conclude with the relation-focused interpretation that target objects are extracted from the relation-related attributes which are enriched at the last-relation token from early layers till middle-late layers and the MLP sublayers are essential in the update of relation representations. Under the guidance of this interpretation, we achieve editing by modifying the MLP sublayer in end of aggregation of relational knowledge (i.e. in the middle-late layer) with the relation representation (i.e. at the last-relation token) while taking subjects into account.
Knowledge Editing from the Relation-focused Perspective
@@
Figure 5: The performance of RETS (purple lines) editing at the last-relation token on different layers (x-axis) compared with ROME (orange line) editing at the last-subject token for 50 prompts. Std deviation is shown in areas.
@@
Our editing experiments are mainly conducted on GPT2-XL and GPT-J for each single factual association. We also evaluate the basic editing performance on Llama-2 (7B). To compare with, we choose the methods for editing each single factual association as our baselines, including Constrained Fine-Tuning (FT+L) (Zhu et al. 2020), the meta-learning method (MEND) (Mitchell et al. 2022) which learns the update of model weights with additional networks, ROME (Meng et al. 2022a) on the subject representation in the early site and the improved precise model editing (PMET) (Li et al. 2024a) which optimizes the parallel MHSA and MLP representations simultaneously.
For evaluation, we conduct experiments on $10,000$ ($2,000$) samples of COUNTERFACT (Meng et al. 2022a) dataset for GPT2-XL (GPT-J), $1,000$ samples of COUNTERFACT for Llama-2 and $10,000$ samples of Zero-shot Relation Extraction (zsRE) (Levy et al. 2017; Meng et al. 2022a) for GPT2-XL. We edit on the $36$-th layer of GPT2-XL, $18$-th layer of GPT-J and $23$-th layer of Llama-2. For COUNTERFACT, we supplement this dataset with unrelated facts of the same subject for each target editing and the corresponding metric Relation Specificity (R-Specificity) to measure the over-generalizing problem. To be specific, given the edited fact <$s_{1},r_{1},o_{1}^{*}$> and the unrelated fact with the same subject <$s_{1},r_{2},o_{2}$>, we test $\mathbb{P}[o_{2}]>\mathbb{P}[o_{1}^{*}]$ as the R-Specificity Score with the input prompt of <$s_{1},r_{2}$>, computed similarly to the existing metrics. We also test the original metrics including the basic Efficacy accuracy Scores to measure the success rate of target editing, the Generalization accuracy score for generalization on the paraphrased statements, the renamed Subject Specificity S-Specificity accuracy score for specificity within neighborhood subjects. The advanced Fluency and Consistency scores measure the quality of generated full texts. Higher scores indicate better performance for all metrics. For zsRE, the metics and evaluation results are displayed in Appendix G. Details for the construction of R-Specificity samples and detailed settings are presented in Appendix F.
@@
Table 3: The evaluation results on COUNERTFACT for GPT2-XL and GPT-J. The significantly failed values for the editing methods on basic criteria are underlined. "Score" shows the average value on the basic criteria: Efficacy, Generalization, S-Specificity and R-Specificity. "SC" stands for the subject constraints on our relation-focused editing. R-Specificity values for raw models are $100.0\%$ since the criterion is constructed according to the top token predictions of the raw models. *PMET is adjusted to accommodate to edit a single layer.
@@
The ablation of the subject constraints for editing COUNTERFACT leads to the $35\%$ decrease on Entity Specificity for GPT2-XL, which indicates the effectiveness of the subject constraints for the relation-focused editing. Despite of the results on Generalization and Entity Specificity where RETS loses about $20\%$ and $10\%$ respectively compared with the subject-centered editing methods, RETS exhibits the most balanced performance with its simple way of combining the subject information into editing, which shows the potential of editing from relation-focused perspective. The trade-off of performance is decided by the editing position (the last-relation token or the last-subject token) as expected. Editing at the last-relation position ensures minimal impact on unrelated relations (i.e. high R-Specificity) but loses much subject information (i.e. low S-Specificity), while the opposite is also true for editing at the last-subject position. The superiority of the relation-focused approach is that the decline of S-Specificity can be constrained by subject constraints whereas the subject-focused approach can hardly attend to
Table 4: An anecdotal example of the correct behavior for RETS and the incorrect behavior for ROME on GPT2-XL. Predictions in red denote unexpectedly changed answers.
the relations. Detailed discussions can be referred to Appendix J.
@@
We test the effectiveness of RETS while editing on different layers and compare it with the behavior of ROME which edits at the last-subject token . Figure 5 plots the performance on four criteria (a,b,c,d) and the average scores (e) across layers. The performance of RETS vibrates before middle layers and become stable after middle-late layers, validating our interpretation of relation-focused knowledge recall that the object is attracted from relation-related attributes which are accumulated before middle-late layers. RETS editing on middle-late layers shows more balanced performance than editing on any layer at the last-subject token where relation information behind hardly propagates to.